# How to Divide: A Set Partitioning Strategy Balancing the Trade-off Between Intra-Subset Correlation and Inter-Subset Gain Mutual Influence in Distributed Attack Detection Scheduling Task

### v1

@@ Abstract
Recently, the efficiency of attack detection in large-scale sensor networks has remained a critical research challenge. Studies have shown that while distributed algorithms offer higher efficiency compared to centralized approaches, they often come at the cost of reduced performance. To strike a balance between detection efficiency and performance in large-scale sensor networks, this paper explores the feasibility of extending existing algorithms to a distributed framework. Starting from the perspective of set partitioning strategies, this study analyzes the key factor that contributes to the performance differences between distributed and centralized algorithms. By examining the gain mutual influence of sensor subsets, an optimal set partitioning strategy is designed to minimize inter-subset mutual influence while enhancing intra-subset correlation. To further reduce the computational cost of gain updates, a suboptimal partitioning strategy based on Grassmann distance is proposed, improving the efficiency of selecting suspicious sensors. Theoretical analysis demonstrates that this approach effectively reduces the computational cost of gain updates while maintaining detection performance. Finally, simulation results validate the effectiveness of the proposed method in enhancing attack detection performance.
Index Terms:
Networks security, large-scale network, attack detection scheduling, set partitioning strategy, security analysis.
I Introduction
@@
With the rapid development of information technology, automation and intelligent systems are profoundly reshaping the way modern society operates. As a key supporting technology for this change, sensor networks have been widely used in various fields, including smart cities, the Internet of Things, and industrial manufacturing [1, 2, 3]. These distributed networks consist of a large number of interconnected sensor nodes that work together to perform environmental monitoring, data collection, and intelligent control, greatly improving the system’s intelligence and autonomous decision-making capabilities [4, 5].
@@
However, as the scale of sensor networks continues to expand, application scenarios are becoming increasingly complex, and security challenges are becoming more prominent, especially in open and dynamic network environments [6]. Adversaries may interfere with the system’s state estimation by means of data tampering and information deception, resulting in system malfunctions or even serious security incidents, bringing significant risks to infrastructure and key businesses [7, 8]. Therefore, in the context of large-scale distributed sensor networks, the development of an efficient and accurate security defense mechanism has emerged as a crucial research direction to ensure system stability and reliability.
@@
Attack detection has long been a core area of research in cybersecurity, encompassing critical domains such as intrusion detection, anomaly detection, and fault diagnosis [9, 10, 11, 12, 13, 14, 15]. Existing methods primarily fall into two categories: rule-based and learning-based approaches. Rule-based approaches rely on predefined security policies and pattern matching techniques, such as anomaly threshold settings and signature detection. For example, Wang et al. [9] extracted topological relationships measured by individual nodes and used reconstruction residuals to pinpoint the location of injection attacks. Suo et al. [10] explored the Grassmann distance of opinion evolution direction and leveraged its latent characteristics to iteratively filter out malicious agents. Li et al. [11] constructed a dynamic relationship between raw data and decrypted results to enable rapid detection of malicious attacks. In the realm of manufacturing system security, Li et al. [12] developed detection rules based on processing procedures and key parameters, identifying network attacks through rule matching. However, rule-based approaches struggle to adapt to novel and complex attacks. In contrast, learning-based approaches leverage data-driven techniques such as statistical analysis, machine learning, and deep learning to detect and recognize attack patterns, offering superior generalization and adaptability [13]. For instance, Zhao et al. [14] designed a data-driven attack detector based on subspace identification, detecting attacks by computing the system’s stable kernel representation. Balta et al. [15] proposed a detection framework based on digital twins, identifying attacks by analyzing controlled transient behaviors. Nevertheless, existing research indicates that these methods still face significant challenges in model training and data dependency [12, 13]. Thus, improving adaptability and computational efficiency while ensuring detection accuracy remains a critical research direction in attack detection.
@@
Beyond the research of attack detection, enhancing the system’s resilience is also crucial for ensuring stability. Secure state estimation aims to accurately recover the state of system even in the presence of attack. This problem has become a key research focus in the field of cyber-physical system security, particularly in control systems and distributed sensor networks [16, 17]. In distributed sensor networks, redundant information is considered essential for ensuring secure state estimation. Even when some sensors are compromised, redundancy allows the system to maintain normal operation [18]. However, identifying compromised sensors efficiently while avoiding the combinatorial explosion associated with NP-hard problems remains a major research challenge [19, 20, 21]. To enhance system resilience, researchers have proposed various secure state estimation strategies. For instance, when core sensors are compromised, Xin et al. [22] introduced an estimation method based on virtual sensors, integrating deep reinforcement learning for online optimization to improve estimation accuracy and reliability. This approach aligns well with the concept of redundant information. Xia et al. [23] proposed a robust distributed Kalman filtering algorithm, which leverages attack detection and robust data fusion strategies to mitigate the impact of malicious network attacks, thereby enhancing the accuracy and stability of distributed state estimation. However, in large-scale distributed sensor networks, computational complexity remains a significant challenge. Balancing security and computational efficiency continues to be a pressing issue that requires further investigation in this field.
@@
For large-scale networks, a sufficient condition to ensure system security is that each sensor can directly exclude potential attacked sensors from the neighbor set. For instance, the ADS algorithm proposed in literature [24] aims to achieve this objective. However, the algorithm is centralized for each node, which means that its efficiency may decrease as the neighborhood cardinality increases. Given that a distributed approach can enhance the efficiency of a centralized algorithm, this paper explores the feasibility of extending the ADS algorithm into a distributed form. Existing studies indicate that distributed algorithms generally lag behind centralized ones in performance, with partitioning strategies considered one of the key factors influencing this gap [25].
@@
To more effectively design partitioning strategies, this paper will extract the inherent characteristics of each sensor. Based on this, the paper aims to optimize the design of partitioning strategies to balance the combined effect of inter-subset benefits and intra-subset correlations. Through optimal partitioning strategy design, the goal of this paper is to narrow the performance gap between distributed and centralized ADS algorithms, achieving a trade-off between efficiency and performance in the suspicious sensor selection task.
@@
II Problem Formulation
II-A System Model
Consider a distributed sensor network that monitors the discrete-time linear time-invariant system state $x(k)$, as described in equation (1)
@@ II-A System Model
$\hat{x}_{i}\left(k+1\right)$ $=$ $A\hat{x}_{i}\left(k\right)+K_{i}\left(k\right)\left(y_{i}\left(k\right)-C_{i}\hat{x}_{i}\left(k\right)\right)$ (3)
-- $\lambda A\sum_{j\in\mathcal{N}_{i}}{\left(\hat{x}_{i}\left(k\right)-\hat{x}_{j}\left(k\right)\right)},$
@@
II-B Attacker Model
At each moment $k$, the attacker selects communication links to launch FDIAs based on the dynamic attack strategy defined in Definition 1.
Definition 1.
(Dynamic Attack Strategy)[24] For the sets of compromised sensors at two consecutive moments, $\mathcal{A}_{k}$ and $\mathcal{A}_{k-1}$, if they differ, i.e., $\mathcal{A}_{k}\neq\mathcal{A}_{k-1}$, we refer to this as a dynamic attack strategy. The difference between the two sets can be calculated as
@@ II-B Attacker Model
  • At any moment $k$, the number of compromised sensors $q_{i}$ in the neighborhood of sensor $i$ does not exceed half the size of the neighborhood, i.e., $q_{i}\leq\lfloor|\mathcal{N}_{i}|/2\rfloor$[26, 21]
@@
II-C Suspicious Sensor Selection Problem
Lemma 1 in literature [24] states that selecting the suspicious sensor set from the set $\mathcal{N}_{i}$ is an NP-hard problem and that the objective function can be transformed into a submodular function as shown in equation (6)
@@ II-C Suspicious Sensor Selection Problem
Problem 1.
[24] The problem of selecting suspicious sensors essentially involves choosing no more than $q_{i}$ sensors to maximize the objective function (6), i.e.,
$\max_{\mathcal{A}_{i,k}\subseteq\mathcal{N}_{i}}f_{k}\left(\mathcal{A}_{i,k}\right),~{}~{}\text{s.t.}\left|\mathcal{A}_{i,k}\right|\leq q_{i},$ (7)
@@ III Main Results
III-A The mutual influence of benefits between subsets
In this section, the neighbor set of each sensor is divided into several subsets, and the ADS algorithm in the literature [24] is deployed on each subset. At this time, the ADS algorithm becomes a distributed case, that is, the D-ADS algorithm.
@@ III-A The mutual influence of benefits between subsets
The method of approximate average partitioning is adopted to ensure that the cardinality of each subset is approximately equal [^1] At this point, the balance between the number of attacked sensors in each subset and the attack intensity can be obtained intuitively. And the results from literature [24] indicate that the D-ADS algorithm guarantees a performance lower bound for each subset consistent with that of the ADS algorithm. The proof process is omitted here. Please refer to the Lemma 2 in APPENDIX A and B.. The cardinality $|\mathcal{N}_{i,g_{i}}|$ of the $g_{i}$-th subset can be expressed as:
@@
However, to ensure this performance lower bound, the gain values of all sensors need to be updated after each selection, and existing literature shows that this will bring huge computational complexity [25]. In fact, at the $l$-th selection at moment $k$, only the sensor $j_{g_{i}^{s},select}^{l_{g_{i}^{s}}}$ in the $g_{i}^{s}$-th subset is selected. To reduce the amount of calculation, this section starts with the gain update method and explores the feasibility of only updating the gains of the remaining elements of the subset where the selected sensor is located.
@@
$\min_{q_{i},j_{q_{i}}}\ \ \sum_{q_{i}=1,q_{i}\neq g_{i}}^{m_{i}}\sum_{j_{q_{i}}\in\mathcal{N}_{i,q_{i}}}E(j_{g_{i}},j_{q_{i}}),$ (11)
@@
III-B The Suboptimal set partitioning strategy based on Grassmann distance
@@
Inspired by the fact that Grassmann distance describes the angular differences in the directions of the vector spanning the subspace, the directional differences between the observable spaces of sensors in the subset are used to evaluate the correlation between sensors. First, the definition of Grassmann distance is introduced:
Definition 3.
(Grassmann Distance)[27] For two subspaces $U$ and $V$, the Grassmann distance $d_{G}(U,V)$ can be calculated by performing singular value decomposition on the bases of these two subspaces, i.e.,
@@
Algorithm 3.1 Sensor Partitioning Strategy Based on Grassmann Distance
@@
According to equation (16), when the indicator function vectors of two sensors are identical, they are fully correlated. In this case, their observable space dimensions must be the same. However, the converse does not necessarily hold. This complete the proof. ∎
@@
Assuming that the diminishing marginal benefit of selecting sensor $j_{s}^{(1)}$ on the gain of $j_{s}^{(2)}$ is ignored, the gain $G_{k,j_{s}^{(2)}}^{(2)}$ of sensor $j_{s}^{(2)}$ is calculated as $G_{k,j_{s}^{(2)}}^{(2)}=f_{k}(\{j_{s}^{(2)}\})$. However, if the influence of selecting element $j_{s}^{(1)}$ on the gain of element $j_{s}^{(2)}$ is considered, the gain $G_{k,j_{s}^{(2)}}^{(2)}$ of sensor $j_{s}^{(2)}$ is calculated as $G_{k,j_{s}^{(2)}}^{(2)}=f_{k}(\mathcal{A}_{i,k}^{(1)}\cup\{j_{s}^{(2)}\})-f_{k}(\mathcal{A}_{i,k}^{(1)})$. The absolute value of the gain error in the two cases is $\Delta_{j_{s}^{(1)},j_{s}^{(2)}}$, which actually indicates the change in the gain of $j_{s}^{(2)}$ caused by the influence of $j_{s}^{(1)}$ on the gain. Therefore, in the two cases, the distribution ratio vectors of the subset where the element $j^{(2)}$ belongs are respectively as shown in equation (18):
$p^{\prime}=\frac{[f_{k}(\{j^{(2)}\}]_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}}{\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})},\quad\quad p^{\prime\prime}=\frac{[f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}]_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}}{\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}})}.$ (18)
@@
$|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|=\frac{|f_{k}(\{j_{s}^{(2)}\}\cdot\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}\Delta_{j_{s}^{(1)},j^{(2)}}-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})\cdot\Delta_{j_{s}^{(1)},j_{s}^{(2)}}|}{(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))}$ (19)
@@
Furthermore, consider the case where there exists attack. Assume that $q_{i,g_{i}}$ sensors in the $g_{i}$-th subset are attacked, while the remaining $|\mathcal{N}_{i,k,g_{i}}|-q_{i,g_{i}}$ sensors are normal.
Divide the equation (19) into two parts: $|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,k,g_{i}}|-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})|$ and $\Delta_{j_{s}^{(1)},j_{max}^{(2)}}/\left((\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))\right)$.
@@
Considering the two parts of equation (19), ideally, $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\rightarrow 0$, because a bounded positive number multiplied by an infinitesimal number tends to 00. In practice, we only need to ensure $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\leq\epsilon$, which is obviously achievable. This also means that based on the proposed set partitioning strategy, only the gains of the remaining sensors in the subset where the selected sensor is located need to be updated, instead of the gains of all the remaining sensors, which will significantly reduce the calculation of the gains. This complete the proof. ∎
IV Simulation Results
In this section, simulation experiments using the vehicle target state estimation are presented to illustrate the effectiveness of the proposed algorithm. Consider the following moving vehicle model[28],

### v2

@@ Abstract
With the large-scale integration of Internet of Things (IoT) into enterprise information management systems, organizations are pursuing digital transformation that hinges on real-time data insights—and yet face escalating security and governance risks. Detecting and responding to threats at scale without impairing system efficiency has therefore become a critical information-management and decision-support challenge for today’s executives. This paper develops a distributed, gain-based anomaly-detection framework tailored to IoT-enabled enterprise systems, underpinned by an optimized sensor-subset partitioning strategy. Starting from the perspective of set partitioning strategies, this study analyzes the key factor that contributes to the performance differences between distributed and centralized algorithms. By examining the gain mutual influence of sensor subsets, an optimal set partitioning strategy is designed to minimize inter-subset mutual influence while enhancing intra-subset correlation. To further reduce the computational cost of gain updates, a suboptimal partitioning strategy based on Grassmann distance is proposed, improving the efficiency of selecting suspicious sensors. Theoretical analysis demonstrates that this approach effectively reduces the computational cost of gain updates while maintaining detection performance. Finally, simulation results validate the effectiveness of the proposed method in enhancing attack detection performance.
keywords:
[…]
1 Introduction
@@
With the rapid development of information technology, automation and intelligent systems are profoundly reshaping the way modern society operates. As a key supporting technology for this change, sensor networks have been widely used in various fields, including smart cities, the Internet of Things, and industrial manufacturing (Heo et al., 2009; Habibzadeh et al., 2017; Hatami et al., 2022; Wang et al., 2020). By enabling comprehensive environmental monitoring, real-time data acquisition, and autonomous control, these networks significantly enhance the intelligence and decision-making capabilities of enterprise information management systems (Wei et al., 2022; Tao et al., 2022; Ju et al., 2023).
@@
However, the proliferation of sensor deployments introduces complex security and governance challenges, as organizational decision makers must balance efficient system operation with the timely detection and response to emerging threats(Xie et al., 2018). Especially in open and dynamic network environments, adversaries may interfere with the system’s state estimation by means of data tampering and information deception, resulting in system malfunctions or even serious security incidents, bringing significant risks to infrastructure and key businesses (Ge et al., 2019; Pang et al., 2021; Smith et al., 2021). Therefore, in the context of large-scale distributed sensor networks, the development of an efficient and accurate security defense mechanism has emerged as a crucial research direction to ensure system stability and reliability.
@@
Attack detection has long been a core area of research in cybersecurity, encompassing critical domains such as intrusion detection, anomaly detection, and fault diagnosis (Wang et al., 2025; Suo et al., 2024a; Li et al., 2023, 2024; Zhang et al., 2021; Zhao et al., 2023; Balta et al., 2023). Existing methods primarily fall into two categories: rule-based and learning-based approaches. Rule-based approaches rely on predefined security policies and pattern matching techniques, such as anomaly threshold settings and signature detection. For example, Wang et al. (2025) extracted topological relationships measured by individual nodes and used reconstruction residuals to pinpoint the location of injection attacks. Suo et al. (2024a) explored the Grassmann distance of opinion evolution direction and leveraged its latent characteristics to iteratively filter out malicious agents. Li et al. (2023) constructed a dynamic relationship between raw data and decrypted results to enable rapid detection of malicious attacks. In the realm of manufacturing system security, Li et al. (2024) developed detection rules based on processing procedures and key parameters, identifying network attacks through rule matching. However, rule-based approaches struggle to adapt to novel and complex attacks. In contrast, learning-based approaches leverage data-driven techniques such as statistical analysis, machine learning, and deep learning to detect and recognize attack patterns, offering superior generalization and adaptability (Zhang et al., 2021). For instance, Zhao et al. (2023) designed a data-driven attack detector based on subspace identification, detecting attacks by computing the system’s stable kernel representation. Balta et al. (2023) proposed a detection framework based on digital twins, identifying attacks by analyzing controlled transient behaviors. To address traffic anomaly detection, Zhang et al. (2022) designed a model consisting of two adversarial sub-networks to learn the data distribution of normal traffic. And based on this, an anomaly detection method based on high anomaly suppression is proposed. Nevertheless, existing research indicates that these methods still face significant challenges in model training and data dependency (Li et al., 2024; Zhang et al., 2021). Thus, improving adaptability and computational efficiency while ensuring detection accuracy remains a critical research direction in attack detection.
@@
Beyond the research of attack detection, enhancing the system’s resilience is also crucial for ensuring stability. Secure state estimation aims to accurately recover the state of system even in the presence of attack. This problem has become a key research focus in the field of cyber-physical system security, particularly in control systems and distributed sensor networks (Ding et al., 2020; Mustafa et al., 2022). In distributed sensor networks, redundant information is considered essential for ensuring secure state estimation. Even when some sensors are compromised, redundancy allows the system to maintain normal operation (Shoukry et al., 2017). However, identifying compromised sensors efficiently while avoiding the combinatorial explosion associated with NP-hard problems remains a major research challenge (Lu and Yang, 2023b; An and Yang, 2022; Lu and Yang, 2023a). To enhance system resilience, researchers have proposed various secure state estimation strategies. For instance, when core sensors are compromised, Xin et al. (2025) introduced an estimation method based on virtual sensors, integrating deep reinforcement learning for online optimization to improve estimation accuracy and reliability. This approach aligns well with the concept of redundant information. Xia and Zhou (2025) proposed a robust distributed Kalman filtering algorithm, which leverages attack detection and robust data fusion strategies to mitigate the impact of malicious network attacks, thereby enhancing the accuracy and stability of distributed state estimation. However, in large-scale distributed sensor networks, computational complexity remains a significant challenge. Balancing security and computational efficiency continues to be a pressing issue that requires further investigation in this field.
@@
For large-scale networks, a sufficient condition to ensure system security is that each sensor can directly exclude potential attacked sensors from the neighbor set. For instance, the ADS algorithm proposed in literature Suo et al. (2024c) aims to achieve this objective. However, the algorithm is centralized for each node, which means that its efficiency may decrease as the neighborhood cardinality increases. Given that a distributed approach can enhance the efficiency of a centralized algorithm, this paper explores the feasibility of extending the ADS algorithm into a distributed form. Existing studies indicate that distributed algorithms generally lag behind centralized ones in performance, with partitioning strategies considered one of the key factors influencing this gap (Mirzasoleiman et al., 2016).
@@
Existing studies have shown that partitioning large-scale sets into subsets based on feature characteristics is an effective way to address the curse of dimensionality (Wang et al., 2023). The key challenge lies in identifying and selecting informative features. To more effectively design partitioning strategies, this paper will extract the inherent characteristics of each sensor. Based on this, the paper aims to optimize the design of partitioning strategies to balance the combined effect of inter-subset benefits and intra-subset correlations. Through optimized partitioning strategy design, the goal of this paper is to narrow the performance gap between distributed and centralized ADS algorithms, achieving a trade-off between efficiency and performance in the suspicious sensor selection task.
@@
2 Problem Formulation
2.1 System Model
IoT systems usually observe production data and system operation status, which usually changes continuously. Therefore, as a theoretical type of article, the following mathematical model is used to describe this process. Consider a distributed sensor network that monitors the discrete-time linear time-invariant system state $x(k)$, as described in equation (1)
@@ 2.1 System Model
$\hat{x}_{i}\left(k+1\right)$ $=$ $A\hat{x}_{i}\left(k\right)+K_{i}\left(k\right)\left(y_{i}\left(k\right)-C_{i}\hat{x}_{i}\left(k\right)\right)-\lambda A\sum_{j\in\mathcal{N}_{i}}{\left(\hat{x}_{i}\left(k\right)-\hat{x}_{j}\left(k\right)\right)},$ (3)
@@
2.2 Attacker Model
To describe the attacker model in IoT, this paper considers a dynamic attack strategy (Suo et al., 2024c), At each moment $k$, the attacker selects communication links to launch FDIAs based on the dynamic attack strategy defined in Definition 1.
Definition 1
(Dynamic Attack Strategy) For the sets of compromised sensors at two consecutive moments, $\mathcal{A}_{k}$ and $\mathcal{A}_{k-1}$, if they differ, i.e., $\mathcal{A}_{k}\neq\mathcal{A}_{k-1}$, we refer to this as a dynamic attack strategy. The difference between the two sets can be calculated as
@@ 2.2 Attacker Model
  • At any moment $k$, the number of compromised sensors $q_{i}$ in the neighborhood of sensor $i$ does not exceed half the size of the neighborhood, i.e., $q_{i}\leq\lfloor|\mathcal{N}_{i}|/2\rfloor$Yang and Lv (2021); Lu and Yang (2023a)
@@
2.3 Suspicious Sensor Selection Problem
Lemma 1 in literature Suo et al. (2024c) states that selecting the suspicious sensor set from the set $\mathcal{N}_{i}$ is an NP-hard problem and that the objective function can be transformed into a submodular function as shown in equation (6)
@@ 2.3 Suspicious Sensor Selection Problem
Problem 1
Suo et al. (2024c) The problem of selecting suspicious sensors essentially involves choosing no more than $q_{i}$ sensors to maximize the objective function (6), i.e.,
$\max_{\mathcal{A}_{i,k}\subseteq\mathcal{N}_{i}}f_{k}\left(\mathcal{A}_{i,k}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\left|\mathcal{A}_{i,k}\right|\leq q_{i},$ (7)
@@ 3 Main Results
3.1 The mutual influence of benefits between subsets
In this section, the neighbor set of each sensor is divided into several subsets, and the ADS algorithm in the literature Suo et al. (2024c) is deployed on each subset. At this time, the ADS algorithm becomes a distributed case, that is, the D-ADS algorithm.
@@ 3.1 The mutual influence of benefits between subsets
The method of approximate average partitioning is adopted to ensure that the cardinality of each subset is approximately equal [^1] At this point, the balance between the number of attacked sensors in each subset and the attack intensity can be obtained intuitively. And the results from literature Suo et al. (2024c) indicate that the D-ADS algorithm guarantees a performance lower bound for each subset consistent with that of the ADS algorithm. The proof process is omitted here. Please refer to the Lemma 2 in APPENDIX A and B.. The cardinality $|\mathcal{N}_{i,g_{i}}|$ of the $g_{i}$-th subset can be expressed as:
@@
However, to ensure this performance lower bound, the gain values of all sensors need to be updated after each selection, and existing literature shows that this will bring huge computational complexity Mirzasoleiman et al. (2016). In fact, at the $l$-th selection at moment $k$, only the sensor $j_{g_{i}^{s},select}^{l_{g_{i}^{s}}}$ in the $g_{i}^{s}$-th subset is selected. To reduce the amount of calculation, this section starts with the gain update method and explores the feasibility of only updating the gains of the remaining elements of the subset where the selected sensor is located.
@@
$\min_{q_{i},j_{q_{i}}}\ \ \sum_{q_{i}=1,q_{i}\neq g_{i}}^{m_{i}}\sum_{j_{q_{i}}\in\mathcal{N}_{i,q_{i}}}E(j_{g_{i}},j_{q_{i}}).$ (11)
At this point, an optimal set partitioning algorithm based on the observable discrimination matrix can be proposed, which can be calculated offline.
@@
3.2 The Suboptimal set partitioning strategy based on Grassmann distance
@@
Inspired by the fact that Grassmann distance describes the angular differences in the directions of the vector spanning the subspace, the directional differences between the observable spaces of sensors in the subset are used to evaluate the correlation between sensors. First, the definition of Grassmann distance is introduced.
Definition 3
(Grassmann Distance)Edelman et al. (1998) For two subspaces $U$ and $V$, the Grassmann distance $d_{G}(U,V)$ can be calculated by performing singular value decomposition on the bases of these two subspaces, i.e.,
@@
Algorithm 3.1 The Set Partitioning Strategy Based on Grassmann Distance
@@
According to equation (16), when the indicator function vectors of two sensors are identical, they are fully correlated. In this case, their observable space dimensions must be the same. However, the converse does not necessarily hold. This complete the proof.
@@
Assuming that the diminishing marginal benefit of selecting sensor $j_{s}^{(1)}$ on the gain of $j_{s}^{(2)}$ is ignored, the gain $G_{k,j_{s}^{(2)}}^{(2)}$ of sensor $j_{s}^{(2)}$ is calculated as $G_{k,j_{s}^{(2)}}^{(2)}=f_{k}(\{j_{s}^{(2)}\})$. However, if the influence of selecting element $j_{s}^{(1)}$ on the gain of element $j_{s}^{(2)}$ is considered, the gain $G_{k,j_{s}^{(2)}}^{(2)}$ of sensor $j_{s}^{(2)}$ is calculated as $G_{k,j_{s}^{(2)}}^{(2)}=f_{k}(\mathcal{A}_{i,k}^{(1)}\cup\{j_{s}^{(2)}\})-f_{k}(\mathcal{A}_{i,k}^{(1)})$. The absolute value of the gain error in the two cases is $\Delta_{j_{s}^{(1)},j_{s}^{(2)}}$, which actually indicates the change in the gain of $j_{s}^{(2)}$ caused by the influence of $j_{s}^{(1)}$ on the gain. Therefore, in the two cases, the distribution ratio vectors of the subset where the element $j^{(2)}$ belongs are respectively as shown below:
$p^{\prime}=\frac{[f_{k}(\{j^{(2)}\}]_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}}{\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})},$ (18)
$p^{\prime\prime}=\frac{[f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}]_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}}{\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}})}.$ (19)
@@
$|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|$ (20)
$=$ $\frac{|f_{k}(\{j_{s}^{(2)}\}\cdot\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}\Delta_{j_{s}^{(1)},j^{(2)}}-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})\cdot\Delta_{j_{s}^{(1)},j_{s}^{(2)}}|}{(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))}$
@@
Furthermore, consider the case where there exists attack. Assume that $q_{i,g_{i}}$ sensors in the $g_{i}$-th subset are attacked, while the remaining $|\mathcal{N}_{i,k,g_{i}}|-q_{i,g_{i}}$ sensors are normal. Then, the equation (20) is divided into two parts: $|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,k,g_{i}}|-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})|$ and $\Delta_{j_{s}^{(1)},j_{max}^{(2)}}/\left((\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))\right)$.
@@
Considering the two parts of equation (20), ideally, $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\rightarrow 0$, because a bounded positive number multiplied by an infinitesimal number tends to 00. In practice, we only need to ensure $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\leq\epsilon$, which is obviously achievable. This also means that based on the proposed set partitioning strategy, only the gains of the remaining sensors in the subset where the selected sensor is located need to be updated, instead of the gains of all the remaining sensors, which will significantly reduce the calculation of the gains. This complete the proof.
4 Simulation Results
In this section, simulation experiments using the vehicle target state estimation are presented to illustrate the effectiveness of the proposed algorithm. Consider the following moving vehicle model(Zhou et al., 2022),